the equation of a line y=mx-1. what is the value of m if the line passes through the point (2,3): the equation of a line y=mx-1. what is the value of m if the line passes through the point (2,3)
Answer by Alan3354(1938)

the equation of a line y=mx-1. what is the value of m if the line passes through the point (2,3)
-----------------
Sub 2 for x and 3 for y
3 = m*2 - 1
2m = 4
m = 2
----
y-y1 = m*(x-x1) (x1,y1) is the point (2,3)
y-3 = 2*(x-2)
y-3 = 2x-4
y = 2x-1 slope-intercept form
2x-y = 1 standard form